OpenStudy (anonymous):

Algebra 2 Word Problem: Set up using 2 variables & 2 equations: "Soybean meal is 16% protein and corn meal is 9% protein. How many pounds of each should be mixed together to get a 350-lb mixture that is 12% protein?"

OpenStudy (anonymous):

How many pounds of protein are there in that 350 pound mix?

OpenStudy (anonymous):

It doesn't state. That is the entire information. Now I need to set up 2 equations using 2 variables for the missing information.

OpenStudy (anonymous):

If the 350 pound mixture is 12% protein, then the number of pounds of protein is 12% of 350. Make sense?

OpenStudy (anonymous):

Using that same reasoning, let S = pounds of soybean meal, and let C = pounds of corn meal; the pounds of protein in the soybean mix is 16% of S and the pounds of protein in the cornmeal mix is 9% of C. See where I'm going with this? @rosimary

OpenStudy (anonymous):

I'm a bit confused. I am suppose to set up 2 equations and then solve by either addition/elimination or substitution method. I am trying to figure out how to write the equations.

OpenStudy (anonymous):

Right, I'm trying to show you how you can both equations in terms of pounds of stuff.

OpenStudy (anonymous):

The first equation is easy, that's the total weight of the mixture. pounds of soybeans + pounds of cornmeal = 350 pounds, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so ... s+c=350

OpenStudy (anonymous):

Right, what I was saying before was to convert the percentages of stuff into pounds of stuff.

OpenStudy (anonymous):

It's important to know what "Soybean meal is 16% protein" means. It's a proportion: pounds of protein : pounds of soybeans :: 16 : 100. Therefore pounds of protein in the soybeans = 0.16*S.

OpenStudy (anonymous):

oh. I was putting 16s ... but it's 0.16s

OpenStudy (anonymous):

You could use 16S, but then you'd have to set it equal to 12*350. It would work out the same, you just need to be careful about consistency.

OpenStudy (anonymous):

Ok. Lets see if i can figure out now

OpenStudy (anonymous):

Im stuck as to how to figure this out

OpenStudy (anonymous):

equation one: total pounds of mixture S + C = 350 equation two: total pounds of protein in mixture 0.16S + 0.09C = 0.12*350

OpenStudy (anonymous):

ok. I did not have eq 2 that way. All I had was .16s+.09c=.12

OpenStudy (anonymous):

why are we multiplying 12 by 350 on 2nd equation

OpenStudy (anonymous):

Because that equation is in terms of the pounds of protein in the mix. The total mix weight is 350 pounds, 12% of which is protein.

OpenStudy (anonymous):

Ok, i think i understand. let me try to figure again

OpenStudy (anonymous):

s = 150 and c =200

OpenStudy (anonymous):

Yes, those are correct.

OpenStudy (anonymous):

Ok. I'm closing this question and opening another. I hate word problems. If you can help I would appreciate it and give you another medal. Thanks